# SofT-GRPO: Surpassing Discrete-Token LLM Reinforcement Learning via Gumbel-Reparameterized Soft-Thinking Policy Optimization

### v1

@@ 4.3 Comparison to butt2025soft
(b) Validation reward curve of different hyper-parameters
@@
Figure 3: Smoothed training or validation curves of ablation studies (the dashed background contains the actual data points). (a) discusses the setting of adding Gumbel noise in SofT-GRPO. (b) discusses the setting of top-p=0.95 and the Gumbel-Softmax temperature $\tau_{g}=0.1$ in SofT-GRPO.
4.4 Comparison in Out-of-Domain Datasets
Besides numerical reasoning, we conduct out-of-domain experiments to evaluate the general reasoning ability of SofT-GRPO. As shown in Table 3, the LLMs fine-tuned with SofT-GRPO on numerical queries can still demonstrate advantages from Pass@1 to Pass@32 on a scientific reasoning benchmark (GPQA Diamond) and two code benchmarks (HumanEval and MBPP).
@@ 4.5 Comparison in the Token Efficiency
In addition to the accuracy, the token efficiency of LLMs is also a crucial metric. In Table 6 and Appendix D.1, we compare the token efficiency of baselines and SofT-GRPO. Compared to the No-Finetune results, fine-tuning with SofT-GRPO will effectively reduce the thinking length. Compared to the discrete-token GRPO, SofT-GRPO will not severely increase the number of tokens. Specifically, there is a clear reduction in the thinking length of the LLaMA-3.2-3B-Instruct model, which demonstrates the effectiveness of SofT-GRPO in saving computational consumption.
4.6 Boosting SofT-GRPO with Majority Voting
As demonstrated in Table 1 and Table 3, SofT-GRPO can lead to clear advantages on Pass@32. To further exploit this advantage, in this subsection, we design to boost SofT-GRPO with majority voting (chen2024more). As shown in Table 4, SofT-GRPO with majority-voting can outperform No-Finetune LLM and LLM fine-tuned with discrete-token GRPO across Major@16 (the accuracy of the most common answer in 16 runs) and Major@32. Results exhibit the potential to strengthen the SofT-GRPO-fine-tuned LLMs for better soft-thinking solvers.
@@ 5 Discussion
5.1 Ablation on Noise Added
In SofT-GRPO, we employ the Gumbel Softmax technique for controllable stochasticity. As discussed in Section 2.2, adding Dirichlet noise to predicted probabilities can be another choice, so we compare the performance of the original SofT-GRPO and two of its variants, adding Dirichlet noise or Gaussian noise to probabilities in training. The reward curve in training is shown in Figure 3(a), the validation curve is shown in Figure 3(b), and the final performance of variants is shown in Table 5. LLMs can not learn a clear refinement from the added Dirichlet noise, and adding the Gaussian noise will cause a poorer initial performance. So, these variants cannot perform as well as integrating the original Gumbel noises.
@@ 5.2 Ablation on Hyper-Parameters
SofT-GRPO sets the top-p as 0.95 and the Gumbel temperature as $\tau_{g}=0.1$. To demonstrate the reason for these settings, in Figure 3(c), we compare the training reward curve of the original SofT-GRPO with two variants (varying top-p to 1.0 or $\tau_{g}$ to 0.25). Results show that both variants will cause a collapse in the training process. As analyzed in Appendix D.3, the two variants will cause a substantial improvement in the KL divergence between $\pi_{\theta}$ and $\pi_{\theta_{ref}}$. So, we attribute this kind of collapse to the fact that fine-tuning under the soft-thinking pattern may lead to inputs outside the pre-trained discrete-token embedding space.
@@ 6 Conclusion
This paper presents a powerful RLVR algorithm, SofT-GRPO, to reinforce LLMs under the soft-thinking reasoning pattern. It integrates controllable stochasticity with Gumbel-Softmax and reinforces the soft-thinking policy with Gumbel reparameterization. The proposed SofT-GRPO can demonstrate better numerical, scientific, and code reasoning ability compared to the conventional discrete-token GRPO on Pass@1, especially Pass@16 and Pass@32. It can also be boosted to a superior solver with majority voting. This article emphasizes the prospects of soft-thinking in LLM reasoning. In the future, we will try building effective tricks over SofT-GRPO and apply it to border fields such as Vision Language Models.
@@ Appendix A Related Work
In this section, we will discuss recent developments on discrete-token RLVR and a similar domain of soft-thinking, the latent reasoning methods.
@@ A.1 RLVR on Discrete-Token CoT Reasoning
The Deepseek-R1-zero (guo2025deepseek) has demonstrated remarkable performance with discrete-token RLVR. And there has been a wide collection of RLVR methods for discrete-token CoT policy optimization, e.g., GRPO (shao2024deepseekmath), Dr. GRPO (liu2025understanding), DAPO (yu2025dapo), and Lite PPO (liu2025part). To further improve these algorithms towards some specific goals, recent works focus on developing RLVR methods for efficient responses (feng2025efficient) or controlling the entropy in CoT generation for better exploration-exploitation balance (hao2025rethinking).
To generate concise discrete-token CoTs, arora2025training modifies the reward function by adding penalties on the generation length. qi2025optimizing and dai2025s turn to formulate the fine-tuning process for concise language CoT as a multi-objective optimization task with the trade-off between token efficiency and accuracy. They truncate the generation of discrete-token CoT at several thinking budgets and optimize the overall performance over them.
RLVR methods with entropy control try to make a better balance between exploration and exploitation. They try to control the entropy in the training process to stimulate the exploration within the groups of CoTs in GRPO wang2025arbitrary.
In the future, we plan to have a comprehensive investigation on whether these modifications can also boost the proposed SofT-GRPO.
@@ A.2 Latent Reasoning
Similar to the soft-thinking pattern, latent reasoning methods pass continuous vectors between LLM steps. These methods fully decouple the reasoning process from explicit natural language (which soft-thinking does not do) and perform inference in the hidden space of the model. Generally, latent CoT methods are often diverse from each other and can be mainly divided into token-wise auto-aggressive methods and auxiliary strategies (chen2025latentcot; tan2025think). Token-wise auto-aggressive methods transform the reasoning process into ’soft thoughts’ with dense latent embeddings (hao2024training) or specialized tokens (e.g., pause (goyal2024pause; zelikman2024quietstar)). These methods focus on transferring the original reasoning policy in the language domain to a latent embedding space, including curriculum learning (e.g., Coconut (hao2024training), LightThinker (zhang2025lightthinker), SIM-COT (wei2025sim)), self-distillation (e.g., CODI (shen2025codi)), and one-shot compression (e.g., CoLaR (tan2025think), SynAdapt (wang2025synadapt)). Auxiliary strategies (e.g., SoftCoT (xu2025softcot)) generate latent embeddings from an auxiliary module and inject them into the frozen main model (cheng2024compressed; xu2025softcot++).
Due to the existing token-wise auto-aggressive methods completely treating the language CoTs as the label, these methods can hardly surpass or even reach the performance level of language LRMs. So empirically, these methods can effectively improve the token efficiency compared to language CoT, but there is a clear performance drop. Auxiliary strategies, instead, can effectively prompt the performance of the original LLM, sacrificing the running efficiency.
@@ Appendix B Prompt of Language Reasoning and Latent Generation
In this part, we show the prompt we adopt for reasoning problems, including the in-domain numerical reasoning, out-of-domain GPQA reasoning, and code reasoning. We inherit the prompts in zhang2025soft for out-of-domain benchmarks. We use the same prompt for both the soft-thinking reasoning pattern and the discrete-token reasoning pattern.
@@
The blue part represents the specific question (query $\boldsymbol{Q}$), the brown part represents the possible test cases provided in the MBPP code reasoning benchmark. We use the code provided in zhang2025soft in the verification process of the responses.
@@ Appendix C Motivation and Theoretical Proof
C.1 Motivation: Mismatch in butt2025soft
butt2025soft adds Gaussian noise on the soft-token inputs as follows:
$\boldsymbol{s}_{t}=\sum_{i=1}^{|\mathcal{T}|}p_{i}\cdot\boldsymbol{e}_{i},\quad\hat{\boldsymbol{s}}_{t}=\boldsymbol{s}_{t}+\mathcal{N}(0,\sigma^{2}Id)$ (13)
@@ C.1 Motivation: Mismatch in butt2025soft
$\exists\,\boldsymbol{p}_{1}\neq\boldsymbol{p}_{2},\quad\text{with}\quad\boldsymbol{E}^{\top}\boldsymbol{p}_{1}=\boldsymbol{E}^{\top}\boldsymbol{p}_{2}$ (16)
@@
Appendix D Supplementary of Experiments

### v2

@@ 4.4 Boosting SofT-GRPO with Majority Voting
(b) Validation accuracy curve of variants on added noises
@@
Figure 3: Smoothed training or validation curves of ablation studies (the dashed background contains the data points). (a) discusses the setting of adding Gumbel noise in SofT-GRPO. (b,c) discusses the setting of top-p=0.95 and the Gumbel-Softmax temperature $\tau_{g}=0.1$.
@@ 4.6 Comparison in the Token Efficiency
In addition to accuracy, token efficiency is an important metric for LLMs. As shown in Appendix E.1, SofT-GRPO reduces thinking length compared to No-Finetune, and does not greatly increase tokens compared to discrete-token GRPO. Notably, SofT-GRPO leads to a clear reduction in thinking length for the LLaMA-3.2-3B-Instruct model, highlighting its effectiveness in reducing computational costs.
@@ 5 Discussion
5.1 Ablation on the Added Noise
In SofT-GRPO, we employ the Gumbel-Softmax technique for controllable stochasticity. As discussed in Section 2.2, adding Dirichlet noise can be another choice, so we compare the performance of the original SofT-GRPO and its two variants, adding Dirichlet noise or Gaussian noise to probabilities $p_{i}$ in training. The reward curve in training is shown in Figure 3(a), the validation curve is shown in Figure 3(b), and the final performance of variants is shown in Table 5. Compared to adding the original Gumbel noises, LLMs do not gain improvements using Dirichlet noise, and adding the Gaussian noise will cause a poorer initial performance.
@@ 5.2 Ablation on Hyper-Parameters
SofT-GRPO sets the top-p as 0.95 and the Gumbel temperature as $\tau_{g}=0.1$. To demonstrate the reason for these settings, in Figure 3(c), we compare the training reward curve of the original SofT-GRPO with two variants (varying top-p to 1.0 or $\tau_{g}$ to 0.25). Results show that both variants will cause a collapse in the training process. As analyzed in Appendix E.3, the two variants will cause a substantial increase in the KL divergence between $\pi_{\theta}$ and $\pi_{\theta_{ref}}$. So, we attribute this kind of collapse to the fact that fine-tuning under the soft-thinking pattern may lead to soft-token inputs outside the pre-trained discrete-token embedding space.
5.3 Visualization: Soft-Thinking after SofT-GRPO
In Appendix H, we provide a visualization of the soft-thinking reasoning path after SofT-GRPO fine-tuning. As shown in Figure 11, high probability paths preserve interpretability, and we also observe new emerged distributions with high mass on a token as well as its antonym (Yeah and No), indicating preservation of multiple reasoning branches.
@@ 6 Conclusion
This paper presents a powerful RLVR algorithm, SofT-GRPO, to reinforce LLMs under the soft-thinking reasoning pattern. It integrates controllable stochasticity with Gumbel-Softmax and updates the soft-thinking policy with Gumbel reparameterization. It can demonstrate superior numerical, scientific, and code reasoning ability compared to the conventional discrete-token GRPO on Pass@1, especially Pass@16 and Pass@32. It can also be boosted into a better solver with majority voting. This work demonstrates the advantages of soft-thinking in LLM reasoning. Future works include extending the method for models such as Vision Language Models.
Impact Statement
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix Contents
  1. 1.
    Related Work ........................................................................................................................................................................A
    1. (a)
      RLVR on Discrete-Token CoT Reasoning
    2. (b)
      Latent Reasoning
    3. (c)
      The Working Mechanism of Soft-thinking
  2. 2.
    Prompt of Language Reasoning and Latent Generation ........................................................................................................................................................................B
  3. 3.
    Motivation and Theoretical Proof ........................................................................................................................................................................C
    1. (a)
      Motivation: Mismatch in Butt et al. (2025) ........................................................................................................................................................................C.1
    2. (b)
      Proof for Theorem 3.1 ........................................................................................................................................................................D
    3. (c)
      Derivation of Eq. (8), Eq. (11) and Eq. (12) ........................................................................................................................................................................C.3
  4. 4.
    Detailed Parameters for Training & Testing ........................................................................................................................................................................D
  5. 5.
    Supplementary of Experiments ........................................................................................................................................................................E
    1. (a)
      Supplementary of Token Efficiency
    2. (b)
      Experiments under Different Temperatures
    3. (c)
      Supplementary Ablation on Hyper-Parameters
    4. (d)
      Pass@K up to K=1024
    5. (e)
      P-values for Significance on Pass@K
  6. 6.
    Discussion & Analysis on Results ........................................................................................................................................................................F
  7. 7.
    Baselines & Datasets & Licenses ........................................................................................................................................................................G
    1. (a)
      Baselines
    2. (b)
      Datasets
    3. (c)
      Inference Framework
    4. (d)
      Licenses
  8. 8.
    Visualization ........................................................................................................................................................................H
@@ Appendix A Related Work
In this section, we will discuss recent developments on discrete-token RLVR, a similar domain of soft-thinking, the latent reasoning methods, and the working mechanism of soft-thinking.
@@ A.1 RLVR on Discrete-Token CoT Reasoning
Deepseek-R1-zero (Guo et al., 2025) has demonstrated remarkable performance with discrete-token RLVR. And there has been a wide collection of RLVR methods for discrete-token CoT policy optimization, such as GRPO (Shao et al., 2024), Dr. GRPO (Liu et al., 2025b), DAPO (Yu et al., 2025), and Lite PPO (Liu et al., 2025c). To further improve these algorithms towards some specific goals, recent works focus on developing RLVR methods for efficient responses (Feng et al., 2025) or controlling the entropy in CoT generation for better exploration-exploitation balance (Hao et al., 2025).
To generate concise discrete-token CoTs, Arora and Zanette (2025) modifies the reward function by adding penalties on the generation length. Qi et al. (2025a) and Dai et al. (2025) turn to formulate the fine-tuning process for concise language CoT as a multi-objective optimization task with the trade-off between token efficiency and accuracy. They truncate the generation of discrete-token CoT at several thinking budgets and optimize the overall performance over them.
RLVR methods with entropy control try to make a better balance between exploration and exploitation. They try to control the entropy in the training process to stimulate the exploration within the groups of CoTs in GRPO (Wang et al., 2025a).
This paper only considers GRPO as the only RLVR pipeline, because we believe these modifications (e.g., considering entropy or DAPO, which proposes to remove the KL divergence term) should be orthogonal to the two major contributions of SofT-GRPO: Sampling with Gumbel-Softmax and credit assignment with Gumbel reparameterization. In the future, we would like to do a comprehensive investigation to check this.
@@ A.2 Latent Reasoning
Similar to the soft-thinking pattern, latent reasoning methods pass continuous vectors between LLM steps. These methods fully decouple the reasoning process from explicit natural language (which soft-thinking does not do) and perform inference in the hidden space of the model. Generally, latent CoT methods are often diverse from each other and can be mainly divided into token-wise auto-regressive methods and auxiliary strategies (Chen et al., 2025; Zhu et al., 2025). Token-wise auto-regressive methods transform the reasoning process into ’soft thoughts’ with dense latent embeddings (Hao et al., 2024) or specialized tokens (e.g., pause (Goyal et al., 2024; Zelikman et al., 2024)). These methods focus on transferring the original reasoning policy in the language domain to a latent embedding space, including curriculum learning (e.g., Coconut (Hao et al., 2024), LightThinker (Zhang et al., 2025a), SIM-COT (Wei et al., 2025)), self-distillation (e.g., CODI (Shen et al., 2025)), and one-shot compression (e.g., CoLaR (Tan et al., 2025), SynAdapt (Wang et al., 2025b), Latent-SFT (Deng et al., 2025)). Auxiliary strategies (e.g., SoftCoT (Xu et al., 2025a)) generate latent embeddings from an auxiliary module and inject them into the frozen main model (Cheng and Van Durme, 2024; Xu et al., 2025b; Su et al., 2025).
Due to the existing token-wise auto-regressive methods completely treating the language CoTs as the label, these methods can hardly surpass or even reach the performance level of reasoning LLMs. So empirically, these methods can effectively improve the token efficiency compared to language CoT, but there is a clear performance drop. Auxiliary strategies, instead, can effectively boost the performance of the original LLM, sacrificing the running efficiency.
@@
A.3 The Working Mechanism of Soft-thinking
As illustrated in (Zhang et al., 2025b), soft-thinking is effective because it transforms the exponential path expansion of reasoning into a tractable and parallelizable process, rigorously approximating full path-summation without sacrificing efficiency or accuracy.
  • Continuous Concept Space: Soft Thinking replaces discrete token choices with “concept tokens,” which are probability-weighted mixtures over all token embeddings. This enables the model to represent more abstract and nuanced concepts beyond individual tokens.
  • Parallel Path Exploration: By retaining the full probability distribution at each reasoning step, Soft Thinking allows the model to implicitly explore multiple reasoning trajectories in parallel, rather than committing early to a single discrete path.
Remarkably, LLMs can directly interpret and process soft tokens without the need for additional fine-tuning. Intuitively, this is because soft tokens are constrained to a well-confined space, specifically the convex hull formed by the model’s embedding vectors.
  • First-Order Linearization for Effectiveness: Moreover, the effectiveness of the soft-thinking paradigm can be understood as a recursive first-order (linear) approximation of the full discrete autoregressive path-sum as follows:
    1. 1.
      For discrete reasoning tokens $\boldsymbol{R}=(r_{1},\ldots,r_{|R|})$, answer $\boldsymbol{A}=(a_{1},\ldots,a_{|A|})$, and LLM policy $\pi_{\theta}$, we have:
      $p(\boldsymbol{A}\mid\boldsymbol{Q})=\sum_{R}\prod_{t=1}^{|\boldsymbol{R}|}\pi_{\theta}(r_{t}\mid[\boldsymbol{Q},(r_{1},...,r_{t-1})])\prod_{t=1}^{|\boldsymbol{A}|}\pi_{\theta}(a_{t}\mid[\boldsymbol{Q},\boldsymbol{R},(a_{1},...,a_{t-1})]).$ (13)
    2. 2.
      Soft-Thinking Tokens: At each reasoning step, define the soft token as the expected embedding:
      $p_{t}\sim\pi_{\theta}(\cdot\mid[\boldsymbol{Q},(\boldsymbol{s}_{1},...,\boldsymbol{s}_{t-1})]),\quad\boldsymbol{s}_{t}=\sum_{i=1}^{|\mathcal{T}|}p_{t,i}\cdot e_{i},$ (14)
      where $p_{t}$ is the token probability distribution, $e_{i}$ is the embedding of token $i$.
    3. 3.
      Then we can recursively replace each sum over discrete tokens by soft tokens. For example, for the first reasoning step:
      $p(\boldsymbol{A}|\boldsymbol{Q})$ $=\sum_{r_{1}}\pi_{\theta}(r_{1}|\boldsymbol{Q})\,p(\boldsymbol{A}|\boldsymbol{Q},r_{1})$ (15)
      $\approx p(\boldsymbol{A}|\boldsymbol{Q},\,\mathbb{E}_{r_{1}}[e_{r_{1}}])$ (16)
      $=p(\boldsymbol{A}|\boldsymbol{Q},\,\boldsymbol{s}_{1}).$ (17)
      Here, $\boldsymbol{s}_{1}=\sum_{i=1}^{|\mathcal{T}|}\pi_{\theta}(r_{1}=i|\boldsymbol{Q})\cdot e_{i}$ is the soft token for the first step.
      For later steps, recursively apply the same procedure:
      $p(\boldsymbol{A}|\boldsymbol{Q},\boldsymbol{s}_{1})$ $=\sum_{r_{2}}\pi_{\theta}(r_{2}|\boldsymbol{Q},\boldsymbol{s}_{1})\,p(\boldsymbol{A}|\boldsymbol{Q},\boldsymbol{s}_{1},r_{2})$ (18)
      $\approx p(\boldsymbol{A}|\boldsymbol{Q},\boldsymbol{s}_{1},\boldsymbol{s}_{2}),$ (19)
      $\text{where}\ \boldsymbol{s}_{2}$ $=\sum_{i=1}^{|\mathcal{T}|}\pi_{\theta}(r_{2}=i|\boldsymbol{Q},\boldsymbol{s}_{1})\cdot e_{i}.$ (20)
    4. 4.
      Recursive Soft-Token Approximation: By recursively applying this linearization at each step, we approximate the original intractable path sum by a deterministic soft-token chain:
      $p(\boldsymbol{A}\mid\boldsymbol{Q})\approx\prod_{t=1}^{|\boldsymbol{S}|}\pi_{\theta}(\boldsymbol{s}_{t}\mid[\boldsymbol{Q},(\boldsymbol{s}_{1},...,\boldsymbol{s}_{t-1})])\prod_{t=1}^{|\boldsymbol{A}|}\pi_{\theta}(a_{t}\mid[\boldsymbol{Q},\boldsymbol{S},(a_{1},...,a_{t-1})]).$ (21)
      where $\boldsymbol{S}=(\boldsymbol{s}_{1},...,\boldsymbol{s}_{|\boldsymbol{S}|})$. Thus, each $\boldsymbol{s}_{t}$ replaces the sum over all possible $r_{t}$, acting as a linear surrogate.
However, it is important to note that not all points in the embedding space (i.e., convex hull of token embeddings) are equally interpretable or accessible to the LLM. For particularly smooth or spread-out distributions (making the approximation in Eq. (15) fail), the expressiveness of soft tokens remains limited. We believe this limitation is closely related to the “collapse” phenomenon discussed in Section 5.2, where the model’s reasoning may degrade or repeat when encountering out-of-distribution soft tokens.
@@ Appendix B Prompt of Language Reasoning and Latent Generation
In this part, we show the prompt we adopt for reasoning problems, including the in-domain numerical reasoning, out-of-domain GPQA reasoning, and code reasoning. We inherit the prompts in Zhang et al. (2025b) for out-of-domain benchmarks. We use the same prompt for both the soft-thinking reasoning pattern and the discrete-token reasoning pattern.
@@
The blue part represents the specific question (query $\boldsymbol{Q}$), the brown part represents the possible test cases provided in the MBPP code reasoning benchmark. We use the code provided in Zhang et al. (2025b) in the verification process of the responses.
@@ Appendix C Motivation and Theoretical Proof
C.1 Motivation: Mismatch in Butt et al. (2025)
Butt et al. (2025) adds Gaussian noise on the soft-token inputs as follows:
$\boldsymbol{s}_{t}=\sum_{i=1}^{|\mathcal{T}|}p_{i}\cdot\boldsymbol{e}_{i},\quad\hat{\boldsymbol{s}}_{t}=\boldsymbol{s}_{t}+\mathcal{N}(0,\sigma^{2}Id).$ (22)
@@
$\exists\,\boldsymbol{p}_{1}\neq\boldsymbol{p}_{2},\quad\text{with}\quad\boldsymbol{E}^{\top}\boldsymbol{p}_{1}=\boldsymbol{E}^{\top}\boldsymbol{p}_{2}.$ (25)
@@
C.3 Derivation of Eq. (8), Eq. (11) and Eq. (12)
The Gumbel-Softmax rollout of SofT-GRPO is given in Eq. (9). On the determinism of the mapping. It is crucial to note that, for fixed logits (or probabilities) $\mathbf{p}_{t}$ and Gumbel noises $\boldsymbol{\epsilon}_{t}$, the resulting soft token $\mathbf{s}_{t}$ is a deterministic function of $(\mathbf{p}_{t},\boldsymbol{\epsilon}_{t})$:
$\mathbf{s}_{t}=\mathrm{SoftEmbed}(\mathbf{p}_{t},\boldsymbol{\epsilon}_{t}),$ (28)
Hence, in our analysis, we always track the generative randomness via $\boldsymbol{\epsilon}_{t}$ (consider $\boldsymbol{\epsilon}_{t}$ as the action in RLVR), and not via $\mathbf{s}_{t}$ directly. Cause every $\boldsymbol{g}^{\prime}$ can map to one deterministic soft-token $\boldsymbol{s}_{t}$, but $\boldsymbol{s}_{t}$ can map to multiple $\boldsymbol{g}^{\prime}$, calculating density on $\boldsymbol{g}^{\prime}$ avoids the ill-posedness of attempting to define sampling densities directly in the soft-token space. So, we consider outputting $\boldsymbol{g}^{\prime}$ as an action and put the generation of $\boldsymbol{s}_{t}$ as a part of the environment transition.
Eq. (11): log-likelihood under the old policy.
During rollout (generated by the old policy $\theta_{\mathrm{old}}$), we sample Gumbel noises $\boldsymbol{\epsilon}_{t}=(\epsilon_{t,1},\dots,\epsilon_{t,|\mathcal{T}|})$. The randomness of $\mathbf{s}_{t}$ is entirely attributable to $\boldsymbol{\epsilon}_{t}$, given $\mathbf{p}_{t}$. The joint density factorizes as
$p(\boldsymbol{\epsilon}_{t})=\prod_{i=1}^{|\mathcal{T}|}f_{\mathrm{Gumbel}}(\epsilon_{t,i}),\qquad f_{\mathrm{Gumbel}}(x)=\exp\!\big(-x-\exp(-x)\big),$
where $f_{\mathrm{Gumbel}}$ represents the probability density function of a Gumbel distribution. Hence, the corresponding log-likelihood is
$\log p(\boldsymbol{\epsilon}_{t})=\sum_{i=1}^{|\mathcal{T}|}\log f_{\mathrm{Gumbel}}(\epsilon_{t,i})=\sum_{i=1}^{|\mathcal{T}|}\Big[-\epsilon_{t,i}-\exp(-\epsilon_{t,i})\Big],$
which gives Eq. (11). In this formulation, the density is over $\boldsymbol{\epsilon}_{t}$, while $\mathbf{s}_{t}$ is merely a deterministic function thereof.
Eq. (12): log-likelihood under the current policy.
For off-policy correction, we reuse the same realized perturbations from rollout. Equivalently, we store the perturbed logits.
$g^{\prime}_{t,i}\triangleq\log p^{\mathrm{old}}_{t,i}+\epsilon_{t,i},$
so that, under a new policy $\theta$ with probabilities $p_{t,i}=\pi_{\theta}(i\mid[Q,(\mathbf{s}_{1},\dots,\mathbf{s}_{t-1})])$, the implied noise that would have generated the same $g^{\prime}_{t,i}$ under the new policy is
$\epsilon_{t,i}=g^{\prime}_{t,i}-\log p_{t,i}.$
Therefore, the log-density of the same realized perturbation under the new policy is
$\sum_{i=1}^{|\mathcal{T}|}\log f_{\mathrm{Gumbel}}(g^{\prime}_{t,i}-\log p_{t,i})=\sum_{i=1}^{|\mathcal{T}|}\Big[-(g^{\prime}_{t,i}-\log p_{t,i})-\exp\!\big(-(g^{\prime}_{t,i}-\log p_{t,i})\big)\Big],$
which is exactly Eq. (12). This preserves consistency in credit assignment and importance sampling, as all stochasticity and probability mass are correctly attributed to $\boldsymbol{\epsilon}_{t}$ rather than $\mathbf{s}_{t}$.
Eq. (8): Off-Policy Ratio
Cause the transformation from $\boldsymbol{g}^{\prime}$ to $\boldsymbol{s}_{t}$ is fixed, the reparameterization process of SofT-GRPO considers the added noise as action. So, using the Eq. (11) and Eq. (12), we can get the final result in Eq. 8 for the soft-thinking part.
$\frac{p(\boldsymbol{g}^{\prime}|[\boldsymbol{Q},(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{t-1})],\theta)}{p(\boldsymbol{g}^{\prime}|[\boldsymbol{Q},(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{t-1})],\theta_{\text{old}})}=$ $\frac{\exp\left(\sum_{i=1}^{|\mathcal{T}|}-(g^{\prime}_{i}-\log p_{i})-\exp(-g^{\prime}_{i}+\log p_{i})\right)}{\exp\left(\sum_{i=1}^{|\mathcal{T}|}-\epsilon_{i}-\exp(-\epsilon_{i})\right)}$ (29)
$=$ $\exp\left(\sum_{i=1}^{|\mathcal{T}|}\big(-(g^{\prime}_{i}-\log p_{i})-\exp(-g^{\prime}_{i}+\log p_{i})\big)-\big(-\epsilon_{i}-\exp(-\epsilon_{i})\big)\right).$
Appendix D Detailed Parameters for Training & Testing
All our experiments are performed on 8 $\times$ H200 GPUs in about 2 to 3 days. Compared to discrete-token GRPO, SofT-GRPO requires $\times$2 to $\times$3 in time, mainly due to the more time consumption for soft-thinking rollout (Zhang et al., 2025b; Wu et al., 2025b). We show the experimental configurations in 6 for No-Finetune, 7 for GRPO, and 8 for SofT-GRPO.
Table 6: Parameters of No-Finetune
Parameter Value
No-Finetune Testing under Discrete-Token Reasoning Paradigm
Maximum response length $32768$ tokens
Sampling temperature 0.6
(top-p, top-k) (0.95, 30)
No-Finetune Testing under Soft-Thinking Reasoning Paradigm
Maximum response length $32768$ tokens
Sampling temperature 0.6
Gumbel-temperature $\tau_{g}$ 0.5
(top-p, top-k) (0.95, 30)
Table 7: Parameters of Discrete-Token GRPO
Parameter Value
Discrete-Token GRPO Training
Maximum response length $8192$ tokens
Sampling temperature 1.0
(top-p, top-k) (1, -1)
Group Size $G$ 8
Learning rate $1\times 10^{-6}$
KL loss coefficient $\beta$ 0.001
Policy clipping parameter $\epsilon$ 0.2
Discrete-Token GRPO Testing under Discrete-Token Reasoning Paradigm
Maximum response length $32768$ tokens
Sampling temperature 0.6
(top-p, top-k) (0.95, 30)
GRPO Testing under Soft-Thinking Reasoning Paradigm
Maximum response length $32768$ tokens
Sampling temperature 0.6
Gumbel-temperature $\tau_{g}$ 0.5
(top-p, top-k) (0.95, 30)
Table 8: Parameters of SofT-GRPO
Parameter Value
SofT-GRPO Training
Maximum response length $8192$ tokens
Sampling temperature 1.0
(top-p, top-k) (0.95, 5)
Group Size $G$ 8
Gumbel-temperature $\tau_{g}$ 0.1
Learning rate $1\times 10^{-6}$
KL loss coefficient $\beta$ 0.001
Policy clipping parameter $\epsilon$ 0.2
SofT-GRPO Testing under Soft-Thinking Reasoning Paradigm
Maximum response length $32768$ tokens
Sampling temperature 1.0
Gumbel-temperature $\tau_{g}$ 0.1
(top-p, top-k) (0.95, 5)
Appendix E Supplementary of Experiments